Copper100:

The following expression has been simplified as follows: (12a^-3b^4)^-2 = 1/144a^-5b^2 =b^2/(144a^5 a. What did the student do wrong? Why do you think the student made this error? b. Solve the problem correctly. ​Make sure to show all work!

UkuleleGirl:

simply: \[\left(12a^{-3}b^4\right)^{-2}=\frac{1}{\left(12a^{-3}b^4\right)^2}\]

UkuleleGirl:

then you simplify more\[\frac{1}{\frac{144b^8}{a^6}}\]

UkuleleGirl:

And apply the rule of fractions. so this is what you intially get here\[=\frac{a^6}{144b^8}\]

UkuleleGirl:

Look back at that work and see what the student did wrong. Do you know what they did wrong?

Copper100:

they added the -2 to -3 and 4

UkuleleGirl:

Yes

Copper100:

ok